Fermi National Laboratory
 

Accelerator Update for October 29, 2007 - October 31, 2007

 

Monday October 29

The day shift began with Operations monitoring an Antiproton Source (Pbar) stack (149.0E10) and a Recycler stash (189.3E10), with the Tevatron (TeV) in a dry squeeze, with the Main Injector (MI) and Recycler conducting studies, and with beam to the SciBooNE / MiniBooNE experiments.

Pbar began stacking at 10:34 AM.

At 10:41 AM, Pbar’s lithium lens tripped off.

Operations put the TeV into shot setup at 10:42 AM.

Booster experts used the 400MeV line flying wires at 11:23 AM to look at beam profiles.

At 11:40 AM, a TeV kicker (C:PAK5) showed a thyratron fault. An expert investigated and said he would need an hour to work on the power supply’s filament.

The Neutron Therapy Facility (NTF) reported at 12:10 PM that they had completed treatment on their last patient.

At 12:21 PM, the TeV sector E1 wet engine tripped off on a system fault. The cryo system technicians at CHL turned it back on. They will investigate further after shot setup.

At 12:49 PM, a water expert began filling the MI RF LCW system.

Operations established store 5661 at 1:11 PM with an initial luminosity of 61.77E30.

At 1:39 PM, Operations reported that the MI-60 LCW system and four Recycler RF stations tripped off during the MI LCW fill. Experts began work on the MI-60 LCW system. The Recycler RF stations will get turned back on after the LCW problems get resolved.

At 2:15 PM, Operations reported that the MI RF LCW system fill was completed.

At 2:37 PM, the 750keV area buncher tripped off. It reset.

Pbar resumed stacking at 3:07 PM.

At 3:09 PM, Operations tried to turn on the Recycler RF stations that had tripped during the MI LCW work. They reported that some amplifiers would not turn on. An expert investigated and fixed the problem.

Operations stashed the Recycler with antiprotons at 4:53 PM. Pbar resumed stacking.

Operations stashed the Recycler at 7:23 PM. Pbar resumed stacking.

TeV experts began a study at 8:05 PM.

Operations stashed the Recycler at 11:26 PM. Pbar resumed stacking.

Tuesday October 30

The midnight shift began with Operations monitoring store 5661 (24.47E30) and the stash (192.9E10), with Pbar stacking (23.3E10), and with Sci /MiniBooNE taking beam.

Operations stashed the Recycler at 2:42 AM. Pbar resumed stacking.

At 4:13 AM, the 750 KeV area buncher tripped off. It reset.

At 5:15 AM, a Booster RF station (BRF4) tripped off. It reset.

At 6:32 AM, a Linac RF station (LRF5) tripped off. It reset.

Operations terminated store 5661 at 7:03 AM.

At 7:21 AM, the 750 KeV area buncher tripped off. It reset.

At 7:27 AM, LRF5 tripped off for the first of three times in 30 minutes. It reset.

TeV kicker experts began their repairs on PAK5 at 7:43 AM. They completed their work 30 minutes later.

At 10:55 AM, BRF17 tripped off. It reset.

TeV kicker experts reported at 12:39 PM that they had found and fixed a loose cable on AAK2.

At 12:54 PM, a Pbar Debuncher RF station (DRF1-1) tripped off. Experts investigated.

Cryo system experts began working on the TeV sector E1 wet engine at 1:09 PM.

A Linac expert worked on the 750 KeV area buncher at 1:38 PM.

A Pbar expert worked on the core momentum PLC at 1:51 PM.

At 3:26 PM, Operations reported that a TeV kicker (PAK3) alarmed on a fail to fire fault. Operators investigated and found that it had fired, but not as fast as PAK2 had fired. Experts investigated.

At 4:27 PM, the 750 KeV area buncher tripped off. It reset.

Operations reported at 6:18 PM that the TeV was at 980 GeV and was going through a wet squeeze.

Operations put the TeV into shot setup at 6:41 PM.

After loading the TeV with antiprotons, the TeV suffered a sector E4 quench at 7:40 PM during a test experts were conducting.

Operations stashed the Recycler at 9:32 PM. Pbar resumed stacking.

Operations put the TeV into shot setup at 11:02 PM.

Wednesday October 31

The midnight shift began with the TeV in shot setup, with Operations monitoring the stack (140.2E10) and the stash (54.9E10), and with Sci /MiniBooNE taking beam.

At 12:24 AM, BRF7 tripped off. It reset.

Operations established store 5668 at 12:53 AM with and initial luminosity of 38.5E30.

Operations stashed the Recycler at 1:06 AM. Pbar resumed stacking.

At 2:18 AM, the 750 KeV area buncher tripped off. It reset.

Operations stashed the Recycler at 5:22 AM. Pbar resumed stacking.

The Plans for Wednesday

The plans for today are to allow TeV experts to move the collision points for CDF and D0, for the TeV to conduct a proton only store, and then to go into a shot setup. The Run Coordinator will allow MI, RR, and Pbar to conduct studies. SY experts will take beam to the Switchyard absorber this morning and then to MT6 this afternoon.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/31/2007    email Fermilab